how do I find side x of a triangle with side 95 angle 45 side 110 assume not a right triangle
Answer by stanbon(75887)

how do I find side x of a triangle with side 95 angle 45 side 110 assume not a right triangle
-------
x is a side the 45 degree angle is between the 95 side and the 110 side
---
So "x" is opposite the 45 Degree angle.
Use the Law of Cosines:
x^2 = 95^2+110^2 -2*95*110*cos(45)
x^2 = 6346.47
x = 79.66
=============
Cheers,
Stan H.
==================